OpenStudy (anonymous):

Choose the correct product of (6x − 2)(6x + 2). 36x2 + 4 36x2 − 24x + 4 36x2 + 24x + 4 36x2 − 4

OpenStudy (anonymous):

you need to open the parenthesis 6x*6x+6x*2-2*6x-2*2=36x^2-4 so it's the fourth answer

OpenStudy (anonymous):

but you can also use a simple formula: (a+b)(a-b)=a^2-b^2